OpenStudy (anonymous):

hi guys i really need help! the sum of three consecutive odd integers is 18 less than five times the middle number. find the three integers. ALGEBRATIC SOLUTION OLNY

OpenStudy (acxbox22):

let x be the first integer x+2 would be the second <<<---- middle number x+4 would be the third x+x+2+x+4=5(x+2)-18 3x+6=5x+10-18 3x+6=5x-8 6=2x-8 14=2x x=7

OpenStudy (acxbox22):

now the first integer is 7 the second one must be 7+2=9 the third one is 7+4=11

OpenStudy (anonymous):

THANK YOU!!! SO MUCH ACXBOX22 your awesome:D

OpenStudy (acxbox22):

no problem :)